OpenStudy (anonymous):

Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 2 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). x = 2 x = 3 x = 5 x = 6

Vocaloid (vocaloid):

to fill in the first row of the table, all you need to do is calculate the value of f(x) at each of the x values to fill in the second row of the table, do the same thing except with g(x) the solution will be the x value where f(x) and g(x) are equal

Vocaloid (vocaloid):

to help you get started, I'll calculate f(1) for you f(x) = -x^2+4x+12 f(1) = -(1)^2+4(1)+12 = -1+4+12 = 15 now do the same thing with the other values

OpenStudy (anonymous):

i got it thanks